# Rate of epidemic spreading on complex networks

### v1

@@ Abstract
One problem with heavy-tailed networks is that the rapid infection of hubs depletes the tail of the degree distribution of the remaining nodes. This implies that, as the epidemic progresses, the degree distribution of the infected nodes does not remain invariant as assumed by our theory. To test this explanation, we study the average degree $\bar{k}_{n}$ of infected individuals in generation $n$. In a steady exponential regime, we expect this quantity to be independent of $n$, see Fig. 4-c. However, in scale-free networks, the average degree of a newly infected rapidly decays, see Fig. 4-d, leading to a non-steady spreading behaviour.
@@
Figure 4: Finite-size effects impact early stages of an epidemic on scale-free networks . Panel a: Relative error between the fit and the Euler-Lotka equation as a function of the exponent $\gamma$ and the finite-size parameter $\eta$ (see Methods). The error is always low, except when the network is scale-free ($2<\gamma\leq 3$, between the dashed lines) and the finite-size parameter $\eta$ is large. Panel b: Error as a function of $\gamma$ and the exponent $\alpha$ of the degree-based clustering coefficient $c_{k}\sim(k-1)^{-\alpha}$ which can be interpreted as the strength of the clustering. Panel c: Average degree of individuals infected at the $n$th generation on a Erdős-Rényi graph of size $N=10^{6}$ and average degree $\overline{k}=4$. The dashed line is a power-law fit $\sim n^{-\delta}$ for $n\in[1,6]$. The fit yields $\delta\approx 0$, consistently with the presence of a clear exponential phase. Panel d: Average degree of individuals infected at the $n$th generation for the Youtube network, which has a diverging second moment ($\gamma\approx 2.2$ (see section VI in SI). The average degree decays as a power law with exponent $\delta\approx 4.48$, suggesting the absence of a clear exponential phase.
@@ Discussion
In this paper, we developed a theory describing the exponential spreading rate of epidemics in populations arranged on a complex network. Our theory relates the spreading rate of an epidemic to the infectiousness and the social structure, while disentangling the impact of their distinct properties. Our approach accurately predicts the exponential rate of epidemic spreading on a wide range of model and real networks, aside from networks with heavy-tailed degree distributions where a clear exponential regime is absent.
When analysing real epidemics in their early stage, a departure from an exponential behaviour in the observed number of cases is often interpreted as a consequence of individual response or early containment measures (see, e.g., [29]). Our theory shows that non-exponential epidemic trajectories are relatively common in heavy-tailed networks, thus providing an alternative explanation for these observations.
We have focused for simplicity on undirected, static networks. The use of a static network is justified when the social structures evolve in a slower way than the time scale of the infection process. If however the time-scales are similar, the epidemic dynamics are more complex [30]. For example, bursty social interactions substantially affect epidemic spreading [31]. In principle, interactions could be also weighted, as duration and intensity of contacts also affect the probability of transmission [32]. The approach developed in this paper is flexible enough to be extended to these cases.
The theory developed in this paper has direct practical consequences for estimating the basic reproduction number $R$ [33]. Ideally, in the early stage of an epidemic, one could use the Euler-Lotka equation to estimate the basic reproduction number. However, this approach is fraught with practical difficulties. First, the curves in Fig. 2f show that a small error in the estimate of the exponential spreading rate $\Lambda$ can lead to a large uncertainty on $R$. Second, it is often difficult to reliably estimate the infectiousness distribution (see, e.g., [34, 35, 36]). The theory presented here permits to directly estimate $R$ from key properties of the interaction networks, such as moments of the degree distribution and clustering. These properties could be estimated, to some extent, based on contact tracing data. As a consequence, such approach could lead to improved estimates of the infectivity function as well.
@@ Methods
We here derive the expression for the reproduction number, Eq. (5). We express the average number of infected individuals at generation $n\geq 1$ as
$Z_{n}=\sum_{k_{1}\dots k_{n}}M_{k_{n}k_{n-1}}M_{k_{n-1}k_{n-2}}\dots M_{k_{2}k_{1}}k_{1}P(k_{1})\,,$ (10)
where the matrix $M_{kk^{\prime}}$ is given by Eq. (3). At the first generation, we obtain $Z_{1}=\bar{k}$. For large $n$, we have $Z_{n}\sim R^{n}$, where $R$ is the leading eigenvalue of $M_{kk^{\prime}}$. For an arbitrary level of clustering and of degree correlation, the leading eigenvalue is not easy to compute exactly. We therefore approximate the eigenvalue by assuming weak clustering and treating the degree correlations as a perturbation of the uncorrelated case.
@@
To further simplify the problem, we neglect the dependence on $k$ of the probability that a node of degree $k^{\prime}$ connected to a node of degree $k$ is part of a triangle. This amounts to assume $(k-1)c_{k}/m_{1}\approx 1$, so that Eq. (13) becomes
@@
In the unperturbed case, the reproduction matrix has a single non-zero eigenvalue $R_{\mathrm{unper}}$ with left and right eigenvectors $u_{k}$ and $v_{k}$, normalised such that $\sum_{k}v_{k}u_{k}=1$:
$R_{\mathrm{unper}}$ $=\frac{\overline{k^{2}}-\overline{k}}{\overline{k}}-m_{1},$ (18)
$u_{k}$ $=\frac{(k-1)(1-c_{k})}{R_{\mathrm{unper}}}$ (19)
$v_{k}$ $=\frac{k}{\overline{k}}P(k).$ (20)
At the order of $\mathcal{O}(r)$, the leading eigenvalue is given by:
$R_{\mathrm{pert}}=R_{\mathrm{unper}}+r\sum_{kk^{\prime}}u_{k}\delta M_{kk^{\prime}}v_{k^{\prime}}+\mathcal{O}(r^{2}).$ (21)
$\sum_{kk^{\prime}}u_{k}\delta M_{kk^{\prime}}v_{k^{\prime}}$ $=\frac{1}{R_{\mathrm{unper}}}\sum_{kk^{\prime}}\left[(k-1)(1-c_{k})(k^{\prime}-1)(1-c(k^{\prime}))P^{(1)}(k|k^{\prime})P(k^{\prime})\frac{k^{\prime}}{\overline{k}}\right]-R_{\mathrm{unper}}.$ (22)
Expanding the terms involving clustering coefficients and evaluating them separately, we obtain
$R_{\mathrm{pert}}=R_{\mathrm{unper}}(1-r)+\frac{r}{R_{\mathrm{unper}}}\left(\frac{\overline{k(k-1)^{2}}}{\overline{k}}-2m_{2}+m_{1}^{2}\right)+\mathcal{O}(r^{2}),$ (23)
where we defined $m_{2}=\overline{k(k-1)^{2}c_{k}}/\overline{k}$. At the first order, the right eigenvector associated to the unique leading eigenvalue is given by:
$v_{k}^{(1)}$ $=v_{k}+\frac{r}{R_{\mathrm{unper}}}\sum_{k^{\prime}}\delta M_{kk^{\prime}}v_{k^{\prime}}$
$\propto\left[1-r+\frac{r}{R_{\mathrm{unper}}}\left((k-1)-m_{1}\right)\right]kP(k)\,.$ (24)
As discussed, this eigenvector, if normalised, represents the relative probability that an infected individual has degree $k$ in the exponential regime.
@@
In this section, we present our extrapolation of the exponential spreading rate $\Lambda$ of an epidemic from simulation data. In networks with heavy tails, finite-size effects alter the course of the epidemic at relatively early times. Likely for this reason, we have observed that functions such as a simple logistic regression fail to fit well the course of the epidemic. In contrast, we have found that the generalised logistic function
$I(t)=\frac{N}{\left(1+Qe^{-\Lambda\nu t}\right)^{1/\nu}},$ (25)
well fits the epidemic trajectories for the large majority of networks that we considered. In Eq. (25), $N$ the size of the network, $Q=(N/I_{0})^{\nu}-1$ and $I_{0}$ the initial number of infected individuals. The free parameters are $\Lambda$, $I_{0}$ and $\alpha$. Although each simulation starts with a single infected individual, we keep $I_{0}$ as a free parameter to better capture the behaviour of the epidemic at the early stages. We fit the logarithm of the generalised logistic function, and discard the first $\overline{k}$ infection events of each simulation, i.e. the times until $I(t)=\overline{k}$. The reason is that these infections should belong to the first generation $Z_{1}$, where the reproduction number is $\overline{k}$ rather than $R$.
@@
$\log I(t)=\log I_{0}+\Lambda t\left(1-\left(\frac{I_{0}}{N}\right)^{\nu}\right)+\mathcal{O}(t^{2}).$ (26)
For $N\rightarrow\infty$, we obtain $I(t)\sim\exp(\Lambda t)$ as expected. Equation (26) leads us to define the finite-size parameter $\eta=\left(I_{0}/N\right)^{\nu}$. For $\nu\to 1$, we recover a classic logistic function, while for $\nu\to 0$ the generalised logistic function tends to a Gompertz curve. For the synthetic and empirical networks that we studied, we have found that the finite-size parameter correlates with the curvature of $\log I(t)$ for small $t$, i.e. the deviation from an exponential behaviour (see section III in SI).
@@
The fits of all epidemic trajectories, together with tables presenting all values of the $\Lambda_{\rm FIT}$ and $\Lambda_{\rm EL}$ are available in section IV in SI.

### v2

@@ Abstract
We aim at understanding why, for certain real-world networks, the predicted spreading rate deviates from the fitted one. We find large errors only for scale-free networks in which the finite-size parameter $\eta$ is also large, see Fig. 4-a. We call a network scale-free if the tail of its degree distribution are fitted by a power law of the form $k^{-\gamma}$, with $\gamma\leq 3$, multiplied by a slowly varying function, see [28] and Section LABEL:SI-sec:tail in SI. In contrast, the error is weakly affected by the strength of clustering, see Fig. 4-b. One issue with heavy-tailed networks is that the rapid infection of hubs depletes the tail of the degree distribution of the remaining nodes. This implies that, as the epidemic progresses, the degree distribution of the infected nodes does not remain invariant as assumed by our theory. To test this explanation, we study the average degree $\bar{k}_{n}$ of infected individuals in generation $n$. In a steady exponential regime, we expect this quantity to be independent of $n$, see Fig. 4-c. However, in scale-free networks, the average degree of infected individuals appears to decay as a power-law of their generation, see Fig. 4-d, leading to a non-steady spreading.
@@
Figure 4: Finite-size effects impact early stages of an epidemic on scale-free networks . Panel a: Relative error between the fit and the Euler-Lotka equation as a function of the exponent $\gamma$ and the finite-size parameter $\eta$ (see Methods). The error is always low, except when the network is scale-free ($2<\gamma\leq 3$, region between the two dashed lines) and the finite-size parameter $\eta$ is large. Panel b: Error as a function of $\gamma$ and the exponent $\alpha$ of the degree-based clustering coefficient $c_{k}\sim(k-1)^{-\alpha}$ which can be interpreted as the strength of the clustering. Panel c: Average degree of individuals infected at the $n$th generation on a Erdős-Rényi graph of size $N=10^{6}$ and average degree $\overline{k}=4$. The dashed line is a power-law fit $\sim n^{-\delta}$ for $n\in[1,6]$. The fit yields $\delta\approx 0$, consistently with a steady exponential spreading. Panel d: Average degree of individuals infected at the $n$th generation for the Youtube network, which is scale-free ($\gamma\approx 2.2$ (see Section LABEL:SI-sec:tail in SI). The average degree decays as a power law with exponent $\delta\approx 4.48$ and is therefore not steady.
@@ Discussion
In this paper, we developed a theory describing the exponential spreading rate of epidemics in populations arranged on a complex network. Our theory relates the spreading rate to the infectiousness and the social structure, while disentangling the impact of their distinct properties. Our approach accurately predicts the spreading rate on a wide range of model and real networks, aside from networks with heavy-tailed degree distributions where the exponential regime is hindered by finite-size effects.
When analysing the early stage of real epidemics, a departure from an exponential behaviour in the observed number of cases is often interpreted as a consequence of individual response or early containment measures (see, e.g., [29]). Our results show that non-exponential epidemic trajectories are common in heavy-tailed networks, thus providing an alternative explanation for these observations.
The theory developed in this paper can be also used to estimate the basic reproduction number $R$ [30]. Ideally, in the early stage of an epidemic, one could use the Euler-Lotka equation to deduce $R$. However, this approach is fraught with practical difficulties. First, the curves in Fig. 2f show that a small error in the estimate of the exponential spreading rate $\Lambda$ can lead to a large uncertainty on $R$. Second, it is often difficult to reliably estimate the infectiousness distribution (see, e.g., [31, 32, 33] for the case of COVID-19). Instead, our theory permits to directly estimate $R$ from structural properties of the interaction networks, such as moments of the degree distribution and clustering. These properties could be inferred, to some extent, based on contact tracing data [34]. As a consequence, such approach could lead to improved estimates of the infectivity function as well.
We have focused for simplicity on static networks, which is justified when the social structures evolve in a slower way than the time scale of the infection process. If however the time-scales are similar, epidemic dynamics are more complex [35]. For example, bursty social interactions substantially affect epidemic spreading [36]. In principle, interactions could be also weighted, as duration and intensity of contacts affect the probability of transmission [37]. The approach developed in this paper is flexible enough to be extended to these scenarios.
@@ Methods
We here derive the expression for the reproduction number, Eq. (5). We express the average number of infected individuals at generation $n\geq 1$ in terms of the reproduction matrix introduced in Eq. (3) by summing over the degree of individuals at each generation:
$Z_{n}=\sum_{k_{1}\dots k_{n}}M_{k_{n}k_{n-1}}M_{k_{n-1}k_{n-2}}\dots M_{k_{2}k_{1}}k_{1}P(k_{1})\,.$ (10)
At the first generation, we obtain $Z_{1}=\bar{k}$. For large $n$, we have $Z_{n}\sim R^{n}$, where $R$ is the leading eigenvalue of $M_{kk^{\prime}}$. For an arbitrary level of clustering and of degree correlation, the leading eigenvalue is not easy to compute exactly. We therefore approximate the eigenvalue by assuming weak clustering and treating the degree correlations as a perturbation of the uncorrelated case.
@@
To further simplify the problem, we neglect the dependence on $k$ of the probability that a node of degree $k^{\prime}$ connected to a node of degree $k$ is part of a triangle. This amounts to assume $(k-1)c_{k}/m_{1}=1$, so that Eq. (13) becomes
@@
At the first order in $r$, we obtain
$R_{\mathrm{pert}}=(R_{0}-m_{1})(1-r)+\frac{r}{R_{0}-m_{1}}\left(\frac{\overline{k(k-1)^{2}}}{\overline{k}}-2m_{2}+m_{1}^{2}\right)+\mathcal{O}(r^{2})\,,$ (18)
see SI, where we defined $m_{2}=\overline{k(k-1)^{2}c_{k}}/\overline{k}$. At this order, the right eigenvector associated to the unique leading eigenvalue is given by:
$v_{k}$ $=\frac{kP(k)}{\bar{k}}+\frac{r}{R_{0}-m_{1}}\sum_{k^{\prime}}\delta M_{kk^{\prime}}\frac{k^{\prime}P(k^{\prime})}{\bar{k}}$
$\propto\left[1-r+\frac{r}{R_{0}-m_{1}}\left((k-1)-m_{1}\right)\right]kP(k)\,.$ (19)
This eigenvector, if normalised, represents the relative probability that an infected individual has degree $k$ in the exponential regime.
@@
We describe our extrapolation of the exponential spreading rate $\Lambda$ of an epidemic from simulation data. In networks with heavy tails, finite-size effects alter the course of the epidemic at relatively early times. Likely for this reason, we have observed that functions such as a simple logistic regression fail to fit well the course of the epidemic. In contrast, we have found that the generalised logistic function
$I(t)=\frac{N}{\left(1+Qe^{-\Lambda_{0}\nu t}\right)^{1/\nu}},$ (20)
well fits the epidemic trajectories for the large majority of networks that we considered. In Eq. (20), $N$ the size of the network and $Q=(N/I_{0})^{\nu}-1$ where $I_{0}$ the initial number of infected individuals. Although each simulation starts with a single infected individual, we consider $I_{0}$ as a free parameter to better capture the behaviour of the epidemic at the early stages. The remaining free parameters are $\Lambda_{0}$ and $\alpha$. We fit the logarithm of the generalised logistic function, and discard the first $\overline{k}$ infection events of each simulation, i.e., the times until $I(t)=\overline{k}$. The reason is that these infections should belong to the first generation $Z_{1}$, in which the average number of secondary infections per individual is $\overline{k}$ rather than $R$.
@@
$\log I(t)=\log I_{0}+\Lambda_{0}t\left(1-\left(\frac{I_{0}}{N}\right)^{\nu}\right)+\mathcal{O}(t^{2}).$ (21)
For $N\rightarrow\infty$, we obtain $I(t)\sim\exp(\Lambda_{0}t)$. Equation (21) leads us to define the finite-size parameter $\eta=\left(I_{0}/N\right)^{\nu}$ and the finite-size growth rate $\Lambda=\Lambda_{0}(1-\eta)$. For $\nu\to 1$, we recover a classic logistic function, while for $\nu\to 0$ the generalised logistic function tends to a Gompertz curve. For the synthetic and real-world networks that we studied, we have found that the finite-size parameter correlates with the curvature of $\log I(t)$ for small $t$, i.e. the deviation from an exponential behaviour (see Section LABEL:SI-sec:eta in SI).
@@
The fits of all epidemic trajectories, together with tables presenting all values of the $\Lambda_{\rm FIT}$ and $\Lambda_{\rm EL}$ are presented in Section LABEL:SI-sec:simulations in SI.
Acknowledgements
We thank C. Castellano and I. Neri for comments on a preliminary version of this manuscript. We thank Joshua Weitz for discussions. We are grateful for the help and support provided by the Scientific Computing and Data Analysis section of Core Facilities at OIST.